OpenStudy (anonymous):

My answer is A Simplify i18 Answers: 1 -1 -i i

OpenStudy (raden):

do u meant i^8 ?

OpenStudy (anonymous):

is that \(i^{18}\) ?

OpenStudy (anonymous):

yes @r0n4ld

OpenStudy (anonymous):

\[i^18=(i^2)^9=1^9=1\]

OpenStudy (anonymous):

yayyyy i was right

OpenStudy (anonymous):

thanks guys you rock my socks

OpenStudy (anonymous):

\[i^{18}=i^2 \cdot i^2 \cdot i^2 \cdot i^2 \cdot i^2 \cdot i^2\cdot i^2\cdot i^2\cdot i^2=(-1)^9=-1\] The answer is B

OpenStudy (anonymous):

actually I did that wrong... it's -1.

OpenStudy (anonymous):

\[i^{18} = (i^2)^9 = (-1)^9 = -1.\]

OpenStudy (anonymous):

really? so i was wrong.

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

Yes. Look at @r0n4ld or my calculation.

OpenStudy (anonymous):

thanks!